how many pounds of coffee worth $10 a pound should be added to 20 pounds of coffee worth $4 a pound to get a mixture worth $5 a pound?
Answer by tommyt3rd(5050)

10x+4(20)=5(x+20)
10x+80=5x+100
5x=20
x=4
4 pounds